# WildFeedback: Aligning LLMs With In-situ User Interactions And Feedback

### v1

@@ Abstract
As large language models (LLMs) continue to advance, aligning these models with human preferences has emerged as a critical challenge. Traditional alignment methods, relying on human or LLM annotated datasets, are limited by their resource-intensive nature, inherent subjectivity, and the risk of feedback loops that amplify model biases. To overcome these limitations, we introduce WildFeedback, a novel framework that leverages real-time, in-situ user interactions to create preference datasets that more accurately reflect authentic human values. WildFeedback operates through a three-step process: feedback signal identification, preference data construction, and user-guided evaluation. We applied this framework to a large corpus of user-LLM conversations, resulting in a rich preference dataset that reflects genuine user preferences. This dataset captures the nuances of user preferences by identifying and classifying feedback signals within natural conversations, thereby enabling the construction of more representative and context-sensitive alignment data. Our extensive experiments demonstrate that LLMs fine-tuned on WildFeedback exhibit significantly improved alignment with user preferences, as evidenced by both traditional benchmarks and our proposed user-guided evaluation. By incorporating real-time feedback from actual users, WildFeedback addresses the scalability, subjectivity, and bias challenges that plague existing approaches, marking a significant step toward developing LLMs that are more responsive to the diverse and evolving needs of their users. In summary, WildFeedback offers a robust, scalable solution for aligning LLMs with true human values, setting a new standard for the development and evaluation of user-centric language models.
@@ 1 Introduction
Large language models (LLMs) have become a cornerstone of modern natural language processing (NLP) applications, powering a wide range of tasks from conversational agents to content generation. Despite their impressive capabilities, a significant challenge remains in aligning these models with human preferences, ensuring that their outputs are not only accurate but also aligned with user expectations and values (Bai et al., 2022a; Ouyang et al., 2022; OpenAI et al., 2024; Dubey et al., 2024). Traditional methods for aligning LLMs typically involve instruction tuning and preference training (Ouyang et al., 2022), where models are fine-tuned on curated datasets that consist of human or LLM annotated preference data (Bai et al., 2022a; Ouyang et al., 2022; Cui et al., 2024). However, these approaches face critical limitations: human annotation is resource-intensive and often subjective, while synthetic data generated by models like GPT-4 (OpenAI et al., 2024) can create feedback loops that reinforce the model’s existing biases rather than capturing the true diversity of human preferences.
In response to these challenges, there has been a growing interest in leveraging in-situ user interactions as a source of feedback for training LLMs (Shi et al., 2022; Lin et al., 2024b; Don-Yehiya et al., 2024). This approach seeks to harness the wealth of authentic, real-time feedback provided by users during their interactions with LLMs, which can offer a more dynamic and accurate reflection of user preferences. By directly incorporating feedback from actual users, rather than relying on pre-defined datasets, it becomes possible to align models more closely with the nuanced and evolving preferences of the broader user population.
@@
In this paper, we introduce WildFeedback, a novel framework designed to align LLMs with in-situ user interactions and feedback. WildFeedback addresses the limitations of existing approaches by constructing preference datasets from real user-LLM conversations, specifically focusing on user feedback that naturally occurs during these interactions. The overview of the framework is shown in Figure 1. Our framework comprises three key components: feedback signal identification, preference data construction, and user-guided evaluation. By applying this framework to WildChat (Zhao et al., 2024), a corpus of over 148,000 multi-turn conversations between users and ChatGPT (OpenAI et al., 2024), we create a preference dataset of 20,281 samples [^1] The dataset will be available soon., capturing a wide range of user preferences.
@@
    Introduction of WildFeedback: We present a novel framework that leverages real-time, in-situ user feedback to construct preference datasets that better reflect actual human values, addressing the scalability and subjectivity issues inherent in human-annotated datasets and the biases in synthetic data.
@@
    Robust Data Construction: We adapt and expand on existing user satisfaction estimation techniques, such as SAT/DSAT rubrics, to identify feedback signals in natural conversations. This enables the creation of a nuanced preference dataset that includes both user preferences and corresponding responses, enhancing the effectiveness of fine-tuning LLMs to better align with user expectations.
@@
    User-Guided Evaluation: We propose a user-guided evaluation methodology that aligns the assessment of model performance with real user preferences, providing a more accurate benchmark for evaluating LLMs’ alignment with human values.
Through extensive experiments, we demonstrate that models fine-tuned on WildFeedback show significant improvements in aligning with user preferences, both in automated benchmarks and in our proposed user-guided evaluation framework. This work represents a step forward in creating more user-centric LLMs, with the potential to enhance user satisfaction across a wide range of applications.
@@ 2 Related Work
Incorporating human feedback has been shown to be an effective strategy to align LLMs with human preferences (Ouyang et al., 2022; Bai et al., 2022a; Dubey et al., 2024). However, relying human annotators to provide human feedback is inefficient and resource-intensive, which makes it hard to scale up. Additionally, human preferences are highly subjective. Preferences of a small set of human annotators could misalign with those of the entire population. Accordingly, some researchers aim to supervise AI models by model themselves (Bai et al., 2022b; Lee et al., 2023; Madaan et al., 2023; Burns et al., 2023; Li et al., 2023a). For instance, Bai et al. (2022b) introduced constitutional AI, in which they prompt LLMs to self-refine their own generations given a set of human-defined constitutions. However, relying on model’ own feedback can create a feedback loop where the model’s outputs increasingly reflect its own biases rather than diverse and authentic human perspectives. Recently, researchers have begun exploring the mining of user preferences from natural human-LLM interactions (Shi et al., 2022; Lin et al., 2024b; Don-Yehiya et al., 2024). These approaches aim to capture authentic, real-time user feedback, which provides a more dynamic and accurate reflection of user preferences. Our work builds on this trend by leveraging in-situ user interactions to create preference datasets that better align with actual human values, addressing the limitations of both synthetic and human-annotated preference datasets.
@@
LLM alignment typically consists of two steps: instruction tuning and preference training. Instruction tuning, or supervised finetuning (SFT), aims to finetune models with a set of instruction-response pairs. Early works incorporated various NLP tasks for instruction tuning, demonstrating that LLMs could generalize well across different tasks (Wang et al., 2022; Chung et al., 2022; Ouyang et al., 2022). Subsequent research focused on constructing instruction data by directly distilling from capable LLMs (Wang et al., 2023; Xu et al., 2023). Researchers later recognized that preference training could further boost model performance across various tasks (Ouyang et al., 2022; Dubey et al., 2024). Unlike SFT, preference training requires not only the instruction and the corresponding desired responses but also the corresponding undesired responses. These responses and their rankings could be either annotated by humans (Bai et al., 2022a) or generated by LLMs (Cui et al., 2024). Beyond general-purpose preference datasets, some datasets focus on specific tasks, such as summarization (Wu et al., 2021), model safety (Ji et al., 2023; Shi et al., 2024), and mathematics (Lightman et al., 2023). However, these approaches often rely on curated datasets that are either manually annotated by human experts or generated by models like GPT-4. While these datasets provide a useful foundation, they may not fully capture the complexity and diversity of real-world user interactions. Our work addresses this gap by introducing a framework that leverages real-time feedback from actual users, allowing for more authentic and context-sensitive alignment of LLMs with true human preferences.
@@ 3 WildFeedback
Existing preference datasets often suffer from a mismatch between actual human preferences and those of the annotators. Synthetic preference datasets, such as UltraFeedback (Cui et al., 2024), rely solely on GPT-4 to generate preferences. However, this approach may not accurately capture real human values or nuanced preferences. Relying on synthetic data can create a feedback loop where the model’s outputs increasingly reflect its own biases rather than diverse and authentic human perspectives. On the other hand, preference datasets annotated by human annotators are difficult to scale due to time and budget constraints (Bai et al., 2022a; Ouyang et al., 2022; Dubey et al., 2024). Moreover, human annotators’ preferences can be highly subjective, often differing significantly from those of real users.
To address these challenges, we introduce WildFeedback, a framework designed to align LLMs with in-situ user interactions and feedback. The framework comprises three steps: (1) feedback signal identification, (2) preference data construction, and (3) user-guided evaluation. The pipeline is illustrated in Figure 1. We apply this framework to WildChat (Zhao et al., 2024), a corpus of real user-ChatGPT conversations, and obtained a preference dataset of 20,281 samples.
@@ 3.1 Feedback Signals Identification
To construct preference data from natural human-LLM interactions, we first need to identify conversations that contain feedback signals. This can be achieved through user satisfaction estimation. In multi-turn conversational sessions, a user may explicitly express their satisfaction (e.g., “thank you”) or dissatisfaction (e.g., “revise it”) in their utterances, or implicitly through their interactions with the agent. Lin et al. (2024b) proposed a framework named SPUR that can automatically learn and identify SAT (satisfaction) and DSAT (dissatisfaction) patterns. SPUR generalizes SAT/DSAT rubrics from conversations with annotated thumb feedback by recursively prompting GPT-4. These rubrics can then be used to score a user’s overall satisfaction or dissatisfaction within a given conversation, allowing us to identify utterances containing feedback signals.
@@
In our setting, we adapt the SAT/DSAT rubrics from Lin et al. (2024b) with minor modifications. In total, we use 9 SAT rubrics and 9 DSAT rubrics. The SAT criteria include gratitude, learning, compliance, praise, personal details, humor, acknowledgment, positive closure, and getting there. The DSAT criteria consist of negative feedback, revision, factual error, unrealistic expectation, no engagement, ignored, lower quality, insufficient detail, and style. Detailed definitions of these rubrics can be found in Table 4 and Table 5. To streamline the process, we input these rubrics into GPT-4  [^2] Unless otherwise specified, in all of our experiments, we use GPT-4o with the gpt-4o-0513 engine. For open-weight models, we use Phi-3-mini-4k-instruct, Mistral-7B-Instruct-v0.3, Meta-Llama-3-8B-Instruct. and prompt it to perform the classification at the utterance level. The complete prompt is available in the Appendix A.1. In total, there are 148,715 multi-turn conversations in the WildChat dataset, with approximately 12.8% of the multi-turn conversations containing feedback signals. Detailed statistics are presented in Table 1.
To ensure the reliability of GPT-4’s classification of SAT/DSAT signals, we conducted a validation process using human expert annotators. We randomly sampled 50 multi-turn conversations, totaling over 500 utterances, and assigned 4 expert annotators to perform the same classification task. Each conversation was annotated by at least 2 annotators, resulting in a final Cohen’s Kappa agreement of $\kappa=0.70$ for SAT and $\kappa=0.54$ for DSAT. After completing the annotations, the annotators reviewed and discussed any disagreements, resolving conflicts to establish a ground truth test set of 50 conversations. Our findings indicate that GPT-4’s ability to identify SAT/DSAT signals shows relatively high agreement with human annotators, achieving a Cohen’s Kappa of $\kappa=0.69$ for SAT and $\kappa=0.50$ for DSAT. A detailed breakdown of GPT-4’s performance and the human annotation process are provided in Appendix B.2.
@@
Table 1: Statistics of the SAT/DSAT conversations and utterances.
# Conv. # Utt. # SAT Conv. # DSAT Conv. # SAT Utt. # DSAT Utt.
148,715 628,467 5,447 13,582 8,186 27,711
3.2 Preference Data Construction
@@
After identifying conversations that contain feedback signals using the SAT/DSAT rubrics, we can construct a preference dataset from these signals. Each sample in this dataset consists of four components: the prompt, user preferences, the preferred response, and the dispreferred response. For conversations containing SAT/DSAT signals, we extract the conversation up to the model response that triggers the SAT/DSAT signals and use this as the prompt for our preference data. Next, we analyze user responses marked by SAT/DSAT signals and ask GPT-4 to summarize user preferences based on these feedback signals (e.g., the user prefers concise and direct answers). By systematically applying the SAT/DSAT rubrics to classify user feedback, we can accurately determine which model responses led to user satisfaction or dissatisfaction.
@@
For the preferred and dispreferred responses, we explore 2 different approaches: expert responses and on-policy responses. Specifically, we use GPT-4 for expert response generation, while Phi 3 (Abdin et al., 2024), Mistral (Jiang et al., 2023), and LLaMA 3 (Dubey et al., 2024) are employed for on-policy response generation. For expert responses, responses that trigger DSAT signals (e.g., requests for revision or expressions of dissatisfaction) in the original conversations are directly used as the dispreferred responses. We then prompt GPT-4 to generate the preferred responses using the summarized user preferences as the system prompt. Similarly, for on-policy responses, both the preferred and dispreferred responses are generated by our policy model. The dispreferred responses are generated directly by the policy model, while the preferred responses are generated using the summarized user preferences as the system prompt. However, recognizing that some user preferences might be harmful in nature (e.g., the user prefers explicit content), we take additional safety precautions. When prompting either on-policy models or GPT-4 to generate the preferred responses, we include an extra system instruction: “the response should be safe.” Some conversations are also filtered out automatically by OpenAI moderation API. This approach allows us to build a robust dataset that trains models to better understand and prioritize responses that align with user preferences. The prompt used for preference data construction is provided in Appendix A.2.
3.3 User-guided Evaluation
In addition to the preference data construction framework, we employ user-guided evaluation to better assess models’ alignment with user preferences. Existing automated benchmarks, such as AlpacaEval (Dubois et al., 2024) and MT-Bench (Zheng et al., 2023b), heavily rely on using LLMs as judges. These benchmarks typically prompt models with a set of queries and then ask LLMs like GPT-4 or Claude (Anthropic, 2023) to provide a score or rank the responses of different models. This approach is problematic because it relies heavily on the internal knowledge of LLMs, which are known to be biased towards longer responses or responses generated by themselves (Liu et al., 2024b; Thakur et al., 2024). Additionally, there is a mismatch between the preferences of LLMs as judges and those of humans, leading to evaluations that do not accurately reflect user preferences. Furthermore, using human annotators to rank model responses base on their subjective experiences is also not ideal, as there can be a mismatch between annotators’ preferences and actual user preferences.
To ensure that our evaluation framework accurately reflects real user preferences, we employ user-guided evaluation. Instead of directly asking annotators to subjectively rank model responses, we should ask them to rank those responses based on real users’ preferences. When using LLMs as evaluators, we provide an instance-level checklist to guide their assessments. In our preference data construction module, we not only construct preference data from user-LLM interactions but also summarize user preferences expressed in natural language. These preferences, based on real users’ textual feedback, can be used to align LLMs’s evaluation more closely with real users’ preferences. Our evaluation framework is adapted from WildBench (Lin et al., 2024a), which has been shown to correlate well with human judgement in ranking model performance as an automatic metric. We employ a pairwise evaluation strategy, where GPT-4 compares two different responses to determine which performs better on a given task, using an instance-level, preference-guided checklist to inform the comparison. This metric allows for straightforward comparisons among models, with easily interpretable win/lose rates as intermediate outcomes. The full prompt can be found in Appendix A.3.
@@ 3.3 User-guided Evaluation
Figure 2: Comparison of in-situ user alignment across datasets generated by different models. “Win/Tie/Lose” represents the percentage of instances where the preferred responses win/tie/lose compared to the dispreferred responses in the WildFeedback dataset, prior to filtering. The comparison is made both with and without providing GPT-4 with summarized user preferences as checklists to guide its evaluation.
3.4 WildFeedback Data
To demonstrate that the generated preferred responses align with actual user preferences, we randomly selected 500 samples from the WildFeedback datasets and performed user-guided evaluation (section 3.3), comparing the preferred and dispreferred responses. As explained in Section section 3.2, there are two versions of WildFeedback: the GPT-4 version and the on-policy version, which differ in whether the responses are generated by GPT-4 or the policy model. As shown in Figure 2, we found that without providing the summarized user preferences as checklists, GPT-4 tends to prefer the dispreferred responses in our dataset, which are the model’s zero-shot generations without guidance from summarized user preferences. However, after providing the preferences as checklists to guide the evaluation, GPT-4’s selections more closely align with real users’ preferences. Additionally, we observed that GPT-4 is significantly more steerable than smaller models: over 70% of its preferred responses align with in-situ user preferences, compared to only about 50% for smaller models. Consequently, for on-policy data, we additionally filter out any data that does not align with user preferences.
We also compare WildFeedback with current open-source datasets in Table 2 333 For UltraFeedback, we refer to the pre-processed, binarized version that was used to train Zephyr (Tunstall et al., 2023).. To the best of our knowledge, WildFeedback is the first multi-turn pairwise preference dataset constructed from real human-LLM interactions. It is also the only dataset derived from in-situ user feedback, unlike existing preference datasets that are annotated by human annotators or LLMs, which often fail to fully capture real users’ preferences. Additionally, although OpenAssistant Conversations (OASST1) (Köpf et al., 2023) also include multi-turn conversations, both its prompts and responses are entirely composed by human annotators, making it less reflective of the genuine dynamics of human-LLM interactions. Overall, WildFeedback outperforms existing datasets in accurately representing authentic human-LLM interactions, making it a more reliable resource for developing and evaluating preference-based models.
Table 2: Statistics of existing preference datasets. The average length refers to the number of tokens. The responses of WildFeedback are either extracted from the original conversations or generated by GPT-4, Mistral, Phi 3, or LLaMA 3.

### v3

@@ Abstract
As large language models (LLMs) continue to advance, aligning these models with human preferences has emerged as a critical challenge. Traditional alignment methods, relying on human or LLM annotated datasets, are limited by their resource-intensive nature, inherent subjectivity, misalignment with real-world user preferences, and the risk of feedback loops that amplify model biases. To overcome these limitations, we introduce WildFeedback, a novel framework that leverages in-situ user feedback during conversations with LLMs to create preference datasets automatically. Given a corpus of multi-turn user-LLM conversation, WildFeedback identifies and classifies user feedback to LLM responses between conversation turns. The user feedback is then used to create examples of preferred and dispreferred responses according to users’ preference. Our experiments demonstrate that LLMs fine-tuned on WildFeedback dataset exhibit significantly improved alignment with user preferences, as evidenced by both traditional benchmarks and our proposed checklist-guided evaluation. By incorporating in-situ feedback from actual users, WildFeedback addresses the scalability, subjectivity, and bias challenges that plague existing approaches, marking a significant step toward developing LLMs that are more responsive to the diverse and evolving needs of their users.
WildFeedback: Aligning LLMs With
In-situ User Interactions And Feedback
Taiwei Shithanks: Corresponding authors: taiweish@usc.edu, wang@tamu.edu, longqi.yang@microsoft.com, jenneville@microsoft.com. The work was done when Taiwei Shi, Zhuoer Wang, Ying-Chun Lin, and Zexue He were interns at Microsoft Corporation.∗†, Zhuoer Wang∗‡, Longqi Yang∗⋄, Ying-Chun Lin, Zexue He, Mengting Wan, Pei Zhou, Sujay Jauhar, Sihao Chen, Shan Xia, Hongfei Zhang Jieyu Zhao, Xiaofeng Xu, Xia Song, Jennifer Neville∗⋄ Microsoft Corporation, Purdue University, Texas A&M University, University of California San Diego, University of Southern California
@@ 1 Introduction
Large language models (LLMs) have become a cornerstone of modern natural language processing (NLP) applications, powering a wide range of tasks from conversational agents to content generation. Despite their strengths, aligning LLMs with human preferences remains a challenge (Bai et al., 2022a; Ouyang et al., 2022; OpenAI et al., 2024; Dubey et al., 2024). Traditional alignment methods involve instruction tuning and preference training on curated human or LLM-annotated datasets (Bai et al., 2022a; Ouyang et al., 2022; Cui et al., 2024). However, these approaches face critical limitations: human annotation is resource-intensive and often subjective, while LLM-generated synthetic data risks reinforcing biases instead of capturing diverse human preferences Gautam and Srinath (2024); Wyllie et al. (2024); Chen et al. (2024); Poddar et al. (2024).
In response, recent work explores in-situ user feedback (e.g., upvotes, downvotes, engagement) for LLM training (Shi et al., 2022; Lin et al., 2024b; Don-Yehiya et al., 2024). This approach harnesses authentic user feedback during interactions with LLMs, offering a more dynamic and accurate reflection of user preferences. Rather than relying on static, costly, and misaligned pre-collected data, this method adapts to evolving user needs. However, existing works are limited in scope, either requiring explicit, structured feedback from users or fine-tuning models directly on responses that trigger explicit user feedback.
@@
In this paper, we introduce WildFeedback, a novel framework designed to align LLMs with in-situ user interactions and feedback. WildFeedback addresses the limitations of existing approaches by constructing preference datasets from real user-LLM conversations, specifically focusing on user feedback that naturally occurs during these interactions. The overview of the framework is shown in Figure 1. Our framework comprises three key components: (1) Feedback signal identification, which detects and classifies user feedback, distinguishing between positive and negative signals to infer user preferences; (2) Preference data construction, which transforms these signals into structured preference datasets; and (3) Checklist-guided evaluation, which systematically assesses model responses using an instance-level checklist derived from extracted user preferences as a rubric. This ensures that model improvements are grounded in real user expectations rather than predefined heuristics. To demonstrate the effectiveness of WildFeedback, we apply it to WildChat (Zhao et al., 2024), a dataset containing over 148,000 multi-turn conversations between users and ChatGPT (OpenAI et al., 2024) (see details of WildChat in Appendix E). This process results in a preference dataset of 20,281 samples [^1] The dataset is available at https://huggingface.co/datasets/microsoft/WildFeedback, providing a rich resource for improving LLM alignment with real-world user preferences.
Through extensive experiments, we demonstrate that models fine-tuned on WildFeedback show significant improvements in aligning with user preferences, both in automated benchmarks and in our proposed checklist-guided evaluation framework. This work represents a step forward in creating more user-centric LLMs, with the potential to enhance user satisfaction across a wide range of applications.
@@
    Introduction of WildFeedback: We present a novel framework that leverages in-situ user feedback to construct preference datasets that better reflect actual human values, addressing the scalability and subjectivity issues inherent in human-annotated datasets and the biases in synthetic data.
@@
    Robust Data Construction: We adapt and expand on existing user satisfaction estimation techniques to identify feedback signals in natural conversations. This enables the creation of a nuanced preference dataset that includes both user preferences and corresponding responses, enhancing the effectiveness of fine-tuning LLMs to better align with user expectations.
@@
    Checklist-Guided Evaluation: We propose a checklist-guided evaluation methodology that aligns the assessment of model performance with real user preferences, providing a more accurate benchmark for evaluating LLMs’ alignment with human values.
@@ 2 Related Work
Incorporating human feedback has been shown to be an effective strategy to align LLMs with human preferences (Ouyang et al., 2022; Bai et al., 2022a; Dubey et al., 2024). However, relying human annotators to provide human feedback is inefficient and resource-intensive, which makes it hard to scale up. Additionally, human preferences are highly subjective. A small set of annotators may not represent broader preferences. Accordingly, some researchers aim to supervise AI models by model themselves (Bai et al., 2022b; Lee et al., 2023; Madaan et al., 2023; Burns et al., 2023; Li et al., 2023a). For instance, Bai et al. (2022b) introduced constitutional AI, in which they prompt LLMs to self-refine their own generations given a set of human-defined constitutions. However, relying on model’s own feedback can create a feedback loop where the model’s outputs increasingly reflect its own biases rather than diverse and authentic human perspectives. Recently, researchers have begun exploring the mining of user preferences from natural human-LLM interactions (Shi et al., 2022; Lin et al., 2024b; Don-Yehiya et al., 2024). These approaches capture real-time user feedback for more accurate preference alignment. Our work builds on this trend by leveraging in-situ user interactions to create preference datasets that better align with actual human values, addressing the limitations of both synthetic and human-annotated preference datasets.
@@
LLM alignment typically consists of two steps: instruction tuning and preference training. Instruction tuning, or supervised finetuning (SFT), aims to finetune models with a set of instruction-response pairs. Early works incorporated various NLP tasks for instruction tuning, demonstrating that LLMs could generalize well across different tasks (Wang et al., 2022; Chung et al., 2022; Ouyang et al., 2022). Subsequent research focused on constructing instruction data by directly distilling from capable LLMs (Wang et al., 2023; Xu et al., 2023). Researchers later recognized that preference training could further boost model performance across various tasks (Ouyang et al., 2022; Dubey et al., 2024). Preference training uses desired and undesired responses, either human-annotated (Bai et al., 2022a) or LLM-generated (Cui et al., 2024). Beyond general-purpose preference datasets, some datasets focus on specific tasks, such as summarization (Wu et al., 2021), model safety (Ji et al., 2023; Shi et al., 2024), and mathematics (Lightman et al., 2023). However, these approaches often rely on curated datasets that are either manually annotated by human experts or generated by models like GPT-4 OpenAI et al. (2024). While these datasets provide a useful foundation, they may not fully capture the complexity and diversity of real-world user interactions. Our work addresses this gap by introducing a framework that leverages real-time feedback from actual users, allowing for more authentic and context-sensitive alignment of LLMs with true human preferences.
@@ 3 WildFeedback
Existing preference datasets often suffer from a mismatch between actual human preferences and those of the annotators Chen et al. (2024); Poddar et al. (2024). Synthetic preference datasets, such as UltraFeedback (Cui et al., 2024), rely solely on GPT-4 to generate rankings and determine which responses are preferred or dispreferred. However, this approach may not accurately capture real human values or nuanced preferences. Relying on synthetic data can create a feedback loop where the model’s outputs increasingly reflect its own biases rather than diverse and authentic human perspectives. On the other hand, preference datasets annotated by human annotators are difficult to scale due to time and budget constraints (Bai et al., 2022a; Ouyang et al., 2022; Dubey et al., 2024). Moreover, human annotators’ preferences can be highly subjective, often differing significantly from those of real users Zhang et al. (2024); Fleisig et al. (2023).
To address these challenges, we introduce WildFeedback, a framework designed to align LLMs with in-situ user interactions and feedback. Unlike previous approaches that rely on synthetic responses, our framework directly learns preferences from real-world users, capturing both explicit and implicit feedback signals. The framework comprises three steps: (1) feedback signal identification, (2) preference data construction, and (3) checklist-guided evaluation. The pipeline is illustrated in Figure 1. We apply this framework to WildChat (Zhao et al., 2024), a corpus of real user-ChatGPT conversations , and obtained the WildFeedback dataset, a preference dataset of 20,281 samples.
@@ 3.1 Feedback Signals Identification
To construct preference data from natural human-LLM interactions, we first identify conversations that contain feedback signals. This can be achieved through user satisfaction estimation. In multi-turn conversational sessions, a user may explicitly express their satisfaction (e.g., “thank you”) or dissatisfaction (e.g., “revise it”) in their utterances. Lin et al. (2024b) proposed a framework named SPUR that can automatically learn and identify SAT (satisfaction) and DSAT (dissatisfaction) patterns. SPUR generalizes SAT/DSAT rubrics from conversations with annotated thumb feedback by recursively prompting GPT-4. These rubrics can then be used to score a user’s overall satisfaction or dissatisfaction, allowing us to identify utterances containing feedback signals.
@@
WildFeedback adapts the SAT/DSAT rubrics from Lin et al. (2024b) with minor modifications. In total, we use 9 SAT and 9 DSAT rubrics. The SAT criteria include gratitude, learning, compliance, praise, personal details, humor, acknowledgment, positive closure, and getting there. The DSAT criteria consist of negative feedback, revision, factual error, unrealistic expectation, no engagement, ignored, lower quality, insufficient detail, and style. Detailed definitions of these rubrics can be found in Table 4 and Table 5. To streamline the process, we input these rubrics into GPT-4  [^2] Unless otherwise specified, in all of our experiments, we use GPT-4o with the gpt-4o-0513 engine. For open-weight models, we use Phi-3-mini-4k-instruct, Qwen2-7B-Instruct, Meta-Llama-3-8B-Instruct. and prompt it to perform the classification at the utterance level. The complete prompt is available in the Appendix A.1. In total, there are 148,715 multi-turn conversations in the WildChat dataset, with approximately 12.8% of the multi-turn conversations containing feedback signals. Detailed statistics are presented in Table 1.
Category SAT DSAT Total
# Conversations 5,447 13,582 148,715
# Utterances 8,186 27,711 628,467
@@
Table 1: Statistics of SAT/DSAT in conversations. A conversation is labeled as SAT/DSAT if it contains at least one SAT/DSAT utterance.
To ensure the reliability of GPT-4’s classification of SAT/DSAT signals, we conducted a validation process using human expert annotators. Our findings indicate that GPT-4’s ability to identify SAT/DSAT signals shows relatively high agreement with human annotations, achieving a Cohen’s Kappa of $\kappa=0.69$ for SAT and $\kappa=0.50$ for DSAT, similar to the human performance. A detailed breakdown of GPT-4’s performance and the human annotation process are provided in Appendix B.2.
3.2 Preference Pair Generation
@@
After identifying conversations that contain feedback signals using the SAT/DSAT rubrics, we can construct semi-synthetic preference pairs. Each preference pair sample consists of four components: the prompt, user preferences, the preferred response, and the dispreferred response. For conversations with SAT/DSAT signals, we first analyze user responses marked by these signals and ask GPT-4 to summarize user preferences based on these feedback signals (e.g., the user prefers concise and direct answers). We then extract the conversation up to the model response that triggers the SAT/DSAT signals and use this as the prompt for our preference data.
@@
For preferred and dispreferred response generation, we explore two different approaches: expert responses and on-policy responses. Specifically, we use GPT-4 for expert response generation, while Phi 3 (Abdin et al., 2024), Qwen 2 (Yang et al., 2024), and LLaMA 3 (Dubey et al., 2024) are employed for on-policy response generation. For expert responses, those that trigger DSAT signals in the original conversations are directly used as dispreferred responses (e.g., response 2 in Fig. 1). We then prompt GPT-4 to generate the preferred responses by using summarized user preferences as the system prompt. For on-policy responses, both preferred and dispreferred responses are generated by the policy model. The dispreferred responses are generated directly, whereas the preferred responses are produced using the summarized user preferences as the system prompt. Furthermore, recognizing that some user preferences may be harmful (e.g., preferences for explicit content), we take extra safety precautions. When prompting either the on-policy models or GPT-4 to generate preferred responses, we include an additional system instruction: “The response should be safe.” Some conversations are also automatically filtered by the OpenAI moderation API. The prompt used for preference pair construction is provided in Appendix A.2.
@@
Figure 2: Comparison of in-situ user alignment across datasets generated by different models. “Win/Tie/Lose” represents the percentage of instances where the preferred responses win/tie/lose compared to the dispreferred responses in the WildFeedback dataset, prior to filtering. The comparison is made both with and without providing GPT-4 with summarized user preferences as checklists to guide its evaluation. With checklists, the preferred responses can be better distinguished.
3.3 Checklist-guided Evaluation
Existing automated benchmarks, such as AlpacaEval (Dubois et al., 2024) and MT-Bench (Zheng et al., 2023b), heavily rely on using LLMs as judges. These benchmarks typically prompt models with a set of queries and then ask LLMs like GPT-4 or Claude (Anthropic, 2023) to provide a score or rank the responses of different models. This approach is problematic because it relies heavily on the internal knowledge of LLMs, which are known to be biased towards longer responses or responses generated by themselves (Liu et al., 2024b; Thakur et al., 2024). Additionally, there is a mismatch between the preferences of LLMs as judges and those of humans, leading to evaluations that do not accurately reflect user preferences. Furthermore, using human annotators to rank model responses based on their subjective experiences is also not ideal, as there can be a mismatch between annotators’ preferences and actual user preferences.
In response, we propose checklist-guided evaluation, a general evaluation framework that more accurately reflects real user preferences. In our preference data construction module, we not only construct preference data from user-LLM interactions but also summarize user preferences expressed in natural language. These preferences, based on real users’ textual feedback, can be used to align LLMs’s evaluation more closely with real users’ preferences. Instead of asking human annotators to directly rank model responses, we should ask them to rank those responses based on real users’ preferences. When using LLMs as evaluators, we can provide an instance-level checklist to guide their assessments. Our evaluation framework is adapted from WildBench (Lin et al., 2024a), which has been shown to correlate well with human judgement in ranking model performance as an automatic metric. We employ a pairwise evaluation strategy, where GPT-4 compares two different responses to determine which performs better on a given task, using an instance-level, preference-guided checklist to inform the comparison. This metric allows for straightforward comparisons among models, with easily interpretable win/lose rates as intermediate outcomes. The full prompt can be found in Appendix A.3.
Similar to feedback signal identification (section 3.1), to ensure the reliability of GPT-4 on checklist-guided evaluation, we conducted a validation process using human expert annotators. We found GPT-4 achieves an human agreement of 57.14%, similar to the human-human agreement of 63.27%. A detailed breakdown of GPT-4’s performance and the human annotation process are provided in Appendix C.